# Thompson Sampling-Based Learning and Control for Unknown Dynamic Systems

### v1

@@ Abstract
Thompson sampling (TS) is an effective method to explore parametric uncertainties and can therefore be used for active learning-based controller design. However, TS relies on finite parametric representations, which limits its applicability to more general spaces, which are more commonly encountered in control system design. To address this issue, this work pro poses a parameterization method for control law learning using reproducing kernel Hilbert spaces and designs a data-driven active learning control approach. Specifically, the proposed method treats the control law as an element in a function space, allowing the design of control laws without imposing restrictions on the system structure or the form of the controller. A TS framework is proposed in this work to explore potential optimal control laws, and the convergence guarantees are further provided for the learning process. Theoretical analysis shows that the proposed method learns the relationship between control laws and closed-loop performance metrics at an exponential rate, and the upper bound of control regret is also derived. Numerical experiments on controlling unknown nonlinear systems validate the effectiveness of the proposed method.
Index Terms:
Learning-based Control, Thompson Sampling, Active Learning Control
@@ I Introduction
In many practical scenarios, the technical challenges or high costs of accurately modeling system dynamics have driven the development of data-driven control, whose performance heavily depends on the quality of the data [1, 2, 3]. To address this issue, active learning control (ALC) improves control performance by actively designing the data acquisition process or enhances data utilization efficiency by actively exploring the most informative sampling points [4, 5, 6]. Thus, ALC has become a powerful tool for enhancing control performance in complex and uncertain systems, particularly in applications such as robotics [7], industrial automation [8], and autonomous vehicles [9].
@@
Among other data-driven control methods, ALC designs data acquisition and utilization processes to improve learning and control effectiveness. Depending on the learning objectives, ALC can be categorized into three types: model-oriented, policy-oriented (control law-oriented), and dual control. Model-oriented methods aim to obtain more accurate or reliable models [10, 11, 12, 13, 14]. In [15], the controller is chosen by maximizing the Fisher information to actively learn the Koopman operator dynamics of robots. Policy-oriented methods focus on exploring the impact of different strategies (control laws) [16, 17]. A reward-free reinforcement learning method was proposed in [17], aiming to explore the relationship between different strategies and reward functions. In the above two types of methods, control performance during the exploration process is often a secondary consideration or even ignored. Dual control, on the other hand, utilizes existing information to not only experiment with the system to learn about its behavior and improve control in the future but also effectively control the system [18, 19, 20]. In dual control, exploration is usually conducted under the premise that certain closed-loop properties of the system are guaranteed, such as stability [18], robustness [19], and regret [20], etc.
As an active sampling method capable of balancing exploration and exploitation, the TS method has gained significant attention in recent years [21, 22, 23]. In the TS method, the weight of sampling an element is proportional to its posterior probability, which can be updated using various distributions, such as Gaussian distributions [24], Bernoulli distributions [25], and Poisson distributions [26].
@@
Following the principles of ALC and the effectiveness of TS, a number of works have been proposed to design active learning control strategies using TS. Leveraging the success of numerous applications, the classical TS method has been explored for stochastic control problems with parametric uncertainty. Early works considered optimization problems where the unknown parameters are selected from a finite discrete set. For example, the work in [27] addressed scenarios where unknown parameters were chosen from a finite discrete set, and the control law is designed based on the learned parameters. More related studies include applications in learning parameterized Markov decision processes [28], regularized linear optimization problems [29], channel selection [30], and multi-agent games and consensus problems [23]. Later, in [31], the gap between discrete and continuous spaces was bridged, where the unknown parameters were assumed to belong to a continuous real space. Expanding on these advancements, the ability to effectively learn unknown parameters in continuous spaces has made the TS method more widely applicable in the field of control. For instance, the authors of [32] considered the control problem of stochastic differential systems with diffusion processes and unknown parameters. They not only analyzed the regret of the learning process but also discussed the stability of the controller, thereby providing a comprehensive framework for such systems. Furthermore, a method for efficiently designing linear-quadratic regulators for linear systems was discussed in [33], which highlights the versatility of TS in addressing classical control problems. Beyond the limitations of finite discrete parameter spaces [27] and continuous real spaces [31], TS methods have the potential for further development in optimizing controller design within functional spaces. This motivates the design of a user-friendly and easily generalizable ALC method in this work.
@@
    A control law learning method is proposed through parameterizing a Hilbert space. By constructing Hilbert spaces and a convex hull structure, the proposed method can effectively represent complex nonlinear control laws, which further provides theoretical support for control law optimization. Furthermore, the upper bound of performance degradation (compared to the potential optimal control law) caused by the proposed parameterization method is analyzed (Theorem 1).
@@
    The control performance of the proposed method is also analyzed. The upper bound of control regret between the proposed method and the potential optimal control law is studied. Specifically, the control regret bound is discussed in terms of the impact of function space parameterization and TS mechanism. Our result (Theorem 3) shows that the control regret consists of a constant part caused by function space parameterization and an exponentially convergent part caused by posterior distribution updates.
@@
The remainder of this work is organized as follows: Section II formulates the considered ALC problem; Section III introduces some useful definitions and lemmas; Section IV presents the main results of this work, including the parameterization of the function space, the learning method based on TS, the learning convergence analysis, and the discussion of the control regret; Section LABEL:sec:NumericalExperiments provides numerical experiments to validate the effectiveness of the proposed methods; and Section V concludes this work and discusses potential future works.
@@
Notation: Random variables are denoted by lowercase letters, and their specific realizations are represented by uppercase letters with subscripts. Sample spaces are indicated by letters like $\mathcal{X,Y}$, and the corresponding $\sigma$-field is denoted using $\Sigma$. For example, let ${\bm{x}}=[x_{1},x_{2},\ldots,x_{n}]^{\rm T}$ be the state of a nonlinear system with samples $X_{1},~{}X_{2},\ldots$, which take values in the space $\mathcal{X}$ with $\sigma$-field $\Sigma_{\bm{x}}$. For a random variable ${\bm{x}}$, the expectation, variance, and covariance are denoted as $\mathbb{E}[{\bm{x}}]$, $\mathbb{V}[{\bm{x}}]$, and $\mathbb{C}[{\bm{x}},{\bm{y}}]$, respectively. The inner product of two vectors ${\bm{x}}$ and ${\bm{y}}$ is denoted as $\langle{\bm{x}},{\bm{y}}\rangle$. The Cartesian product of sets $\mathcal{G}_{1},\mathcal{G}_{2}$ is represented as $\mathcal{G}_{1}\times\mathcal{G}_{2}$. Moreover, we define $\underline{\bm{x}}\otimes\overline{\bm{x}}:=\{{\bm{x}}=[{x}_{1},\ldots,{x}_{n}]^{\rm T}|{x}_{i}\in\{\underline{x}_{i},\overline{x}_{i}\}\}$. In this work, $n$-dimensional real spaces are denoted as $\mathbb{R}^{n}$, $\mathbb{N}$ is used to represent the set of natural numbers, and $\mathbb{N}_{n}:=\{1,2,\ldots,n\}$. Let $[a,b]:=\{a,\ldots,b\}$ for $a,b\in\mathbb{N}$ be a shorthand notation for a continuous set of natural numbers. Similarly, let ${\bm{x}}_{[a,b]}:=\{x_{a},\ldots,x_{b}\}$ be a shorthand notation for a set consisting of items of the vector ${\bm{x}}=[x_{a},\ldots,x_{b}]^{\rm T}$. The $\mathcal{L}_{2}$-norm of a function is denoted as $\|\cdot\|_{\mathcal{L}_{2}}$, and the $\mathcal{L}_{2}$-norm of a vector-valued function ${\bm{g}}=[g_{1},\ldots,g_{m}]^{\rm T}$ is defined as $\|{\bm{g}}\|_{\mathcal{L}_{2}}^{2}:=\sum\limits_{i=1}^{m}\|g_{i}\|^{2}_{\mathcal{L}_{2}}$.
@@ II Problem Formulation
where ${\bm{x}}(k)\in\mathcal{X}\subset\mathbb{R}^{n}$ represents the state, ${\bm{v}}(k)$ denotes unknown disturbances, ${\bm{u}}(k)\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, and ${\bm{f}}(\cdot)=[f_{0}(\cdot),\ldots,f_{n}(\cdot)]^{\rm T}$ are unknown system dynamics. In this work, the control input is generated by a controller ${\bm{g}}(\cdot)$, i.e., ${\bm{u}}(k)={\bm{g}}({\bm{x}}(k))=[g_{1}({\bm{x}}(k)),\ldots,g_{m}({\bm{x}}(k))]^{\rm T}\in\mathbb{R}^{m}$.
The control objective of this work is to design a controller ${\bm{g}}(\cdot)$ for the given unknown system (1) and a specific task, such that a cost function $J$ is minimized. In practice, the cost can be computed based on the system states and inputs over time instants $\{0,\ldots,K\}$. A widely used example is the quadratic cost function $J=\sum\limits_{k=0}^{K}({\bm{x}}(k)^{\rm T}{Q}{\bm{x}}(k)+{\bm{u}}(k)^{\rm T}{R}{\bm{u}}(k))$, where ${Q}$ and ${R}$ are positive definite matrices. Another commonly discussed example is the exponential cost function (also referred to as the risk-sensitive cost function [34]) $J=\sum\limits_{k=0}^{K}e^{-\alpha k}({\bm{x}}(k)^{\rm T}{Q}{\bm{x}}(k)+{\bm{u}}(k)^{\rm T}{R}{\bm{u}}(k))$, where $\alpha>0$ is a user-defined parameter. In this work, we assume that the cost function can also be computed from the system states and input data, as illustrated in the aforementioned examples.
@@
There exists a set $\bar{\mathcal{X}}$ satisfying:
@@
$|J({\bm{g}}_{1})-J({\bm{g}}_{2})|\leq L_{J}\|{\bm{g}}_{1}-{\bm{g}}_{2}\|_{\mathcal{L}_{2}},{\bm{g}}_{1},{\bm{g}}_{1}\in\check{\mathcal{G}},$ (2)
and $J({\bm{g}})$ is bounded as $J({\bm{g}})\geq\underline{J_{m}},{\bm{g}}\in\check{\mathcal{G}}$.
Remark 1.
Assumption 1 can be easily satisfied by constructing an appropriate $\bar{\mathcal{X}}$. For example, for any set $\mathcal{X}$, the set $\bar{\mathcal{X}}$ can be defined as: $\bar{\mathcal{X}}=\{{\bm{x}}|\|{\bm{x}}\|_{\infty}\leq\max\limits_{\bar{\bm{x}}\in\mathcal{X}}\|\bar{\bm{x}}\|_{\infty}\}$. The purpose of Assumption 1 is to simplify the structure of the set $\mathcal{X}$ to facilitate the design of the control law. Assumption 2 can also be satisfied when considering bounded sets $\mathcal{X},\mathcal{U}$, which aligns with practical scenarios because of the bounded energy restrictions in practical systems.
@@ IV-A Parameterization of Controllers
For a vector $\tilde{\bm{x}}\in\mathcal{X}$ and a set ${\bm{w}}\subseteq[1,n]$, we denote a vector ${\bm{d}}=[d_{1},\ldots,d_{n}]=:({\bm{x}},\tilde{\bm{x}})_{\bm{w}}$ as follows:
$d_{i}=\left\{\begin{array}[]{ll}x_{i},{}&{\text{if}}~{}i\in{\bm{w}},\\ \tilde{x}_{i},{}&{\text{if}}~{}i\notin{\bm{w}}.\end{array}\right.$ (10)
@@
$\mathcal{G}_{i}:=\Bigg{\{}{g}\Bigg{|}$ ${g}=\sum\limits_{{\bm{w}}\subseteq[1,n]}\alpha({\bm{w}})\Gamma g^{(i)}_{{\bm{w}},\tilde{\bm{x}}},\alpha({\bm{w}})\geq 0,$
$\sum\limits_{{\bm{w}}\subseteq[1,n]}\alpha({\bm{w}})=1,\Bigg{\}},~{}i\in\{1,\ldots,m\},$ (11)
@@
    $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}({\bm{x}})=\sum\limits_{\varpi\subseteq w}(-1)^{|{\bm{w}}|-|{\bm{\varpi}}|}g_{i}(({\bm{x}},\tilde{\bm{x}})_{\bm{\varpi}})$;
@@
with $|{\bm{w}}|$ representing the number of elements in ${\bm{w}}$. The set $\mathcal{G}_{i}$ can be regarded as the convex hull of multiple points in the function space $\check{\mathcal{G}}_{i}$ composed of basis functions $\{g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}\}$.
The parameterization algorithm for the set $\mathcal{G}_{i}$ is summarized in Algorithm 1. The algorithm is organized as a recursive function that computes the basis functions $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}$ based on the initial control law $\check{g}_{i}(\cdot)$ and the constant $\tilde{\bm{x}}$. The algorithm starts with the empty set ${\bm{w}}=\emptyset$, where it directly returns the value of the initial control law at $\tilde{\bm{x}}$. For other cases, it recursively computes the basis functions by summing over all proper subsets of ${\bm{w}}$, ensuring that the contribution of different combinations of variables can be captured by the constructed basis functions.
Algorithm 1 Parameterization Algorithm for $\mathcal{G}_{i}$.
1:  Function: Output=Func($\check{g}_{i}$,${\bm{w}}$,$\tilde{\bm{x}}$);
2:  if ${\bm{w}}=\emptyset$ then
3:     $g^{(i)}_{\emptyset,\tilde{\bm{x}}}=\check{g}_{i}(\tilde{\bm{x}})$;
4:     Output=$g^{(i)}_{\emptyset,\tilde{\bm{x}}}$;
5:  else
6:     $g^{(i)}=0$;
7:     for each ${\bm{\varpi}}\subsetneq{\bm{w}}$ do
8:        $g^{(i)}=g^{(i)}$+Func($\check{g}_{i}$,${\bm{\varpi}}$,$\tilde{\bm{x}}$)
9:     end for
10:     Output=$\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{\bm{w}})-g^{(i)}$;
11:  end if
@@
In this work, we assum that the $\mathcal{L}_{2}$-norm of the considered base functions are bounded, i.e., $\mathcal{G}_{i}\subset\mathcal{G}^{*}$ with $\mathcal{G}^{*}$ defined as
$\mathcal{G}^{*}=\left\{g|\|g\|_{\mathcal{L}_{2}}\leq M_{g}\right\}.$ (13)
@@
For the considered system (1), there exists an optimal controller ${\bm{g}}^{*}=[g_{1}^{*},\ldots,g_{m}^{*}]^{\rm T},~{}g_{i}^{*}\in\mathcal{G}^{*}$ that minimizes the cost function $J({\bm{g}})$.

### v2

@@ Abstract
Thompson sampling (TS) is a Bayesian randomized exploration strategy that samples options (e.g., system parameters or control laws) from the current posterior and then applies the selected option that is optimal for a task, thereby balancing exploration and exploitation; this makes TS effective for active learning-based controller design. However, TS relies on finite parametric representations, which limits its applicability to more general spaces, which are more commonly encountered in control system design. To address this issue, this work proposes a parameterization method for control law learning using reproducing kernel Hilbert spaces and designs a data-driven active learning control approach. Specifically, the proposed method treats the control law as an element in a function space, allowing the design of control laws without imposing restrictions on the system structure or the form of the controller. A TS framework is proposed in this work to reduce control costs through online exploration and exploitation, and the convergence guarantees are further provided for the learning process. Theoretical analysis shows that the proposed method learns the relationship between control laws and closed-loop performance metrics at an exponential rate, and the upper bound of control regret is also derived. Furthermore, the closed-loop stability of the proposed learning framework is analyzed. Numerical experiments on controlling unknown nonlinear systems validate the effectiveness of the proposed method.
@@ I Introduction
In many practical scenarios, such as robotics [11], autonomous vehicles[35], and industrial processes [26], the effectiveness of data-driven control critically depends on the availability of high-quality, informative data that sufficiently captures the system’s operational modes and uncertainties. However, obtaining such representative datasets is challenging in practice due to safety, cost, and operational constraints [12, 13, 33, 30]. To address this issue, active learning control (ALC) improves control performance by actively designing the data acquisition process or enhances data utilization efficiency by actively exploring the most informative sampling points [4, 19, 14]. Thus, ALC has become a powerful tool for enhancing control performance in complex and uncertain systems, particularly in applications such as robotics [38], industrial automation [15], and autonomous vehicles [46].
@@
Among other data-driven control methods, ALC designs data acquisition and utilization processes to improve learning and control effectiveness. Depending on the learning objectives, ALC can be categorized into three types: model-based control, policy-based control, and dual control. Model-oriented methods aim to obtain accurate or reliable models [29, 8, 21, 32, 36]. In [3], the controller is chosen by maximizing the Fisher information to actively learn the Koopman operator dynamics of robots. Policy-oriented methods focus on exploring the impact of different strategies (control laws) [28, 10]. A reward-free reinforcement learning method was proposed in [10], aiming to explore the relationship between different strategies and reward functions. In the above two types of methods, control performance during the exploration process is often a secondary consideration or even ignored. Dual control, on the other hand, utilizes existing information to not only experiment with the system to learn about its behavior and improve control in the future but also effectively control the system [40, 42, 24]. In dual control, exploration is usually conducted under the premise that certain closed-loop properties of the system are guaranteed, such as stability [40], robustness [42], and regret [24], etc.
To balance exploration and exploitation during active sampling, various sampling methods for uncertain scenarios have been proposed, including the Upper Confidence Bound (UCB) [16] and Thompson Sampling (TS) [27, 2, 18]. In the UCB method, the policy is always chosen as the one corresponding to the upper confidence bound of the reward function [9]. In the TS method, the reward corresponding to each option is randomly generated according to the posterior distribution, and the option with the highest reward is selected. The TS can naturally incorporate prior knowledge, effectively handle nonstationary processes (such as time-varying system dynamics), and has been shown to have stronger adaptability and higher convergence speed compared to UCB [44]. Therefore, in this work, the TS framework is chosen to design the data-driven active learning control strategy.
@@
Following the principles of ALC and the effectiveness of TS, a number of works have been proposed to design active learning control strategies using TS. Leveraging the success of numerous applications, the classical TS method has been explored for stochastic control problems with parametric uncertainty. Early works considered optimization problems where the unknown system parameters are selected from a finite discrete set. For example, the work in [23] addressed scenarios where unknown parameters were chosen from a finite discrete set, and the control law is designed based on the learned parameters. More related studies include applications in learning parameterized Markov decision processes [17], regularized linear optimization problems [1], channel selection [25], and multi-agent games and consensus problems [18]. Later, in [5], the gap between discrete and continuous spaces was bridged, where the unknown parameters were assumed to belong to a continuous real space. Expanding on these advancements, the ability to effectively learn unknown parameters in continuous spaces has made the TS method more widely applicable in the field of control. For instance, the authors of [37] considered the control problem of stochastic differential systems with diffusion processes and unknown parameters. They not only analyzed the regret of the learning process but also discussed the stability of the controller, thereby providing a comprehensive framework for such systems. Furthermore, a method for efficiently designing linear-quadratic regulators for linear systems was discussed in [22], which highlights the versatility of TS in addressing classical control problems. Beyond the limitations of finite discrete parameter spaces [23] and continuous real spaces [5], TS methods have the potential for further development in optimizing controller design within functional spaces. This motivates the design of a user-friendly and easily generalizable ALC method in this work.
@@
    A control law learning method is proposed through parameterization of a Hilbert space. By constructing Hilbert spaces and a convex hull structure, the proposed method can effectively represent complex nonlinear control laws, which further provides theoretical support for control law optimization. Furthermore, the upper bound of performance degradation (compared to the potential optimal control law) caused by the proposed parameterization method is analyzed (Theorem 1).
@@
    The control performance of the proposed method is also analyzed. The upper bound of control regret between the proposed method and the optimal control law is studied, and the closed-loop stability is also analyzed. Specifically, we discuss the control regret bound (Theorem 3) and closed-loop stability (Theorem 4) with respect to the impact of function space parameterization and the TS mechanism. Our results show that the control regret comprises a constant part due to function space parameterization and an exponentially decaying part resulting from posterior distribution updates. Furthermore, the state of the closed-loop system is mean-square bounded.
@@
The remainder of this work is organized as follows: Section II formulates the considered ALC problem; Section III introduces some useful definitions and lemmas; Section IV presents the main results of this work, including the parameterization of the function space, the learning method based on TS, the learning convergence analysis, and the discussion of the control regret; Section V provides several extensions and discussions, including reward functions with nonstationary distributions, closed-loop stability, and computational complexity; Section VI provides numerical experiments to validate the effectiveness of the proposed methods; and Section VII concludes this work and discusses potential future works.
@@
Notation: Random variables are denoted by lowercase letters, and their specific realizations are represented by uppercase letters with subscripts. Sample spaces are indicated by letters like $\mathcal{X,Y}$, and the corresponding $\sigma$-field is denoted using $\Sigma$. For example, let ${\bm{x}}=[x_{1},x_{2},\ldots,x_{n}]^{\rm T}$ be the state of a nonlinear system with samples $X_{1},~X_{2},\ldots$, which take values in the space $\mathcal{X}$ with $\sigma$-field $\Sigma_{\bm{x}}$. For a random variable ${\bm{x}}\in\mathcal{X}$, the expectation, variance, and covariance are denoted as $\mathbb{E}[{\bm{x}}]$, $\mathbb{V}[{\bm{x}}]$, and $\mathbb{C}[{\bm{x}}_{1},{\bm{x}}_{2}],~{\bm{x}}_{1},{\bm{x}}_{2}\in\mathcal{X}$, respectively. The inner product of two vectors $\bm{x}$} and $\bm{y}$} is denoted as $\langle{\bm{x}},{\bm{y}}\rangle$. The Cartesian product of sets $\mathcal{G}_{1},\mathcal{G}_{2}$ is represented as $\mathcal{G}_{1}\times\mathcal{G}_{2}$. Moreover, we define $\underline{\bm{x}}\otimes\overline{\bm{x}}:=\{{\bm{x}}=[{x}_{1},\ldots,{x}_{n}]^{\rm T}|{x}_{i}\in\{\underline{x}_{i},\overline{x}_{i}\}\}$. In this work, $n$-dimensional real spaces are denoted as $\mathbb{R}^{n}$, $\mathbb{N}$ is used to represent the set of natural numbers, and $\mathbb{N}_{n}:=\{1,2,\ldots,n\}$. Let $a,b]:=\{a,\ldots,b\}$ for $a,b\in\mathbb{N}$ be a shorthand notation for a continuous set of natural numbers. Similarly, let ${\bm{x}}_{[a,b]}:=\{x_{a},\ldots,x_{b}\}$ be a shorthand notation for a set consisting of items of the vector ${\bm{x}}=[x_{a},\ldots,x_{b}]^{\rm T}$. The $\mathcal{L}_{2}$-norm of a function is denoted as $\|\cdot\|_{\mathcal{L}_{2}}$, and the $\mathcal{L}_{2}$-norm of a vector-valued function ${\bm{g}}=[g_{1},\ldots,g_{m}]^{\rm T}$ is defined as $\|{\bm{g}}\|_{\mathcal{L}_{2}}^{2}:=\sum\limits_{i=1}^{m}\|g_{i}\|^{2}_{\mathcal{L}_{2}}$. Moreover, ‘a.s.’ is used to denote ‘almost surely’.
@@ II Problem Formulation
where ${\bm{x}}(k)\in\mathcal{X}\subset\mathbb{R}^{n}$ represents the state, ${\bm{v}}(k)\in\mathcal{V}\subset\mathbb{R}^{n_{v}}$ denotes unknown disturbances, ${\bm{u}}(k)\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, and ${\bm{f}}(\cdot)=[f_{0}(\cdot),\ldots,f_{n}(\cdot)]^{\rm T}$ are unknown system dynamics. In this work, the control input is generated by a state-feedback controller ${\bm{g}}(\cdot)$, i.e., ${\bm{u}}(k)={\bm{g}}({\bm{x}}(k))=[g_{1}({\bm{x}}(k)),\ldots,g_{m}({\bm{x}}(k))]^{\rm T}\in\mathbb{R}^{m}$.
The control objective of this work is to design a controller ${\bm{g}}(\cdot)$ for the given unknown system (1) and a specific task, such that a cost function $J$ is minimized. In practice, the cost can be computed based on the system states and inputs over time instants $\{0,\ldots,K\}$. A widely used example is the quadratic cost function
$J\!=\!\sum\limits_{k=0}^{K}({\bm{x}}(k)^{\rm T}{Q}{\bm{x}}(k)\!+\!{\bm{u}}(k)^{\rm T}{R}{\bm{u}}(k)),$ (2)
where Q{Q} and R{R} are positive definite matrices. Another widely used example is the risk-sensitive cost function [43]
$J\!=\!\frac{1}{\alpha_{\rm risk}}\!\log\mathbb{E}\!\left[\exp\!\left(\sum\limits_{k=0}^{K}\alpha_{\rm risk}({\bm{x}}(\!k\!)^{\rm T}{Q}{\bm{x}}(\!k\!)\!+\!{\bm{u}}(\!k\!)^{\rm T}{R}{\bm{u}}(\!k\!))\!\right)\!\right]$ (3)
where $\alpha_{\rm risk}>0$ quantifies the degree of risk sensitivity. A higher $\alpha_{\rm risk}$ places greater emphasis on minimizing high-cost scenarios, making the controller more risk-averse, while a lower $\alpha_{\rm risk}$ results in a more risk-neutral behavior.
@@
There exists a convex set $\bar{\mathcal{X}}$ satisfying:
@@
Remark 1.
Assumption 1 introduces a convex envelope $\bar{\mathcal{X}}$ of the feasible state set $\mathcal{X}$ that serves as the design domain for constructing the controller class and for subsequent analysis. In practice, $\bar{\mathcal{X}}$ is chosen by enlarging the known or anticipated operating region to include safety/modeling margins; it does not restrict the actual closed-loop trajectories and can be tuned to balance conservatism and tractability. The condition $\overline{\bm{x}}\otimes\underline{\bm{x}}\subset\bar{\mathcal{X}}$ simply encodes that convex combinations of admissible states remain in $\bar{\mathcal{X}}$. Specifically, $\bar{\mathcal{X}}$ can be set as the minimal axis-aligned bounding box $\{\bm{x}:\|\bm{x}\|_{\infty}\leq\sup_{\bm{z}\in\mathcal{X}}\|\bm{z}\|_{\infty}\}$. Furthermore, when $\mathcal{X}=\mathbb{R}^{n}$, the choice $\bar{\mathcal{X}}=\mathbb{R}^{n}$ can be adopted. In this way, it is ensured that $\mathcal{X}\subseteq\bar{\mathcal{X}}$ and that convex combinations of admissible states remain in $\bar{\mathcal{X}}$.
@@
$|J({\bm{g}}_{1})-J({\bm{g}}_{2})|\leq L_{J}\|{\bm{g}}_{1}-{\bm{g}}_{2}\|_{\mathcal{L}_{2}},~{\bm{g}}_{1},{\bm{g}}_{2}\in\check{\mathcal{G}},$ (4)
and $J({\bm{g}})$ is bounded as $J({\bm{g}})\geq{\bar{J}_{\rm min}},{\bm{g}}\in\check{\mathcal{G}}$ with $\bar{J}_{\rm min}$} being a positive constant.
@@ III-A Neighborhood
III-B Thompson Sampling
TS is a sampling strategy for selecting among multiple options with uncertain rewards. A simple example of TS is provided as follows. Consider a scenario where $N_{\rm dice}$ biased dice are available; in each round, we can choose one of these dice to roll, and the number on the dice is our reward. Our objective is to maximize the reward. Although these dice may appear identical, differences in their mass distribution lead to unknown reward distributions. Therefore, when making a decision (choosing which dice to roll), we need to simultaneously consider two goals: how to gather more information about each dice’s reward distribution (exploration), and how to maximize the reward in the current round (exploitation).
TS provides an analytically convenient strategy to address such problems. The specific procedure is:
  • Step 1: Initialize a probability distribution (such as a uniform distribution) for the reward of each dice;
  • Step 2: For each of the $N_{\rm dice}$ dice, draw one sample from the current probability distribution of each dice, resulting in $N_{\rm dice}$ reward samples;
  • Step 3: Select the dice with the highest sampled reward in Step 2 and roll it;
  • Step 4: Update the posterior distribution of the selected dice based on the observed reward, and repeat from Step 2.
In contrast to the above scenario, the objective of this work is to select a particular controller from a set of candidate control laws. The selected controller is evaluated through the closed-loop system dynamics. By online active sampling, we aim to identify control laws that yield higher rewards (i.e., lower costs).
@@ IV-A Parameterization of Controllers
For $\tilde{\bm{x}}\in\mathcal{X}$, $\forall{\bm{x}}\in\mathcal{X}$, and an index set ${\bm{w}}\subseteq[1,n]$, we denote a vector ${\bm{d}}=[d_{1},\ldots,d_{n}]=:({\bm{x}},\tilde{\bm{x}})_{\bm{w}}$ as follows:
$d_{i}=\left\{\begin{array}[]{ll}x_{i},&{\text{if}}~i\in{\bm{w}},\\ \tilde{x}_{i},&{\text{if}}~i\notin{\bm{w}},\end{array}\right$. (12)
where $\bm{w}$} is a subset of the set $1,n]=\{1,\ldots,n\}$, including both the empty set and the full set. Therefore, there are $2^{n}$ possible $\bm{w}$}, which are used to construct the basis functions in the following context.
@@
$\mathcal{G}_{i}:=\Bigg\{{g}\Bigg|$ ${g}=\sum\limits_{{\bm{w}}\subseteq[1,n]}\alpha_{\bm{w}}\Gamma g^{(i)}_{{\bm{w}},\tilde{\bm{x}}},~\alpha_{\bm{w}}\geq 0,$
$\sum\limits_{{\bm{w}}\subseteq[1,n]}\alpha_{\bm{w}}=1,\Bigg\},~i\in\{1,\ldots,m\},$ (13)
@@
    $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}({\bm{x}})=\sum\limits_{\varpi\subseteq w}(-1)^{|{\bm{w}}|-|{\bm{\varpi}}|}\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{\bm{\varpi}})$;
@@
with $|{\bm{w}}|$ representing the number of elements in $\bm{w}$}.
An example of constructing the function space $\mathcal{G}_{i}$ is provided below.
Example 1.
For an initial controller $\check{g}_{i}({\bm{x}})=x_{1}+x_{2}+x_{1}x_{2}$ and a state $\tilde{\bm{x}}=[1~2]^{\rm T}$, the index set $\bm{w}$} satisfies
${\bm{w}}\in\{{\bm{w}}_{1},{\bm{w}}_{2},{\bm{w}}_{3},{\bm{w}}_{4}\}=\{\emptyset,\{1\},\{2\},\{1,2\}\}.$ (14)
Then the basis function corresponding to ${\bm{w}}_{1}=\emptyset$ can be obtained as $g^{(i)}_{\emptyset,\tilde{\bm{x}}}\overset{\rm i)}{=}\check{g}_{i}(\tilde{\bm{x}})=5$, based on which the basis functions for ${\bm{w}}_{2}$ and ${\bm{w}}_{3}$ can be calculated as
$g^{(i)}_{{\bm{w}}_{2},\tilde{\bm{x}}}\overset{\rm i)}{=}\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{{\bm{w}}_{2}})\!-\!g^{(i)}_{\emptyset,\tilde{\bm{x}}}\overset{(\!1\!0\!)}{=}\check{g}_{i}([x_{1}~2]^{\rm T})-g^{(i)}_{\emptyset,\tilde{\bm{x}}}=3x_{1}-3,$
$g^{(i)}_{{\bm{w}}_{3},\tilde{\bm{x}}}\overset{\rm i)}{=}\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{{\bm{w}}_{3}})\!-\!g^{(i)}_{\emptyset,\tilde{\bm{x}}}\overset{(\!1\!0\!)}{=}\check{g}_{i}([1~x_{2}]^{\rm T})-g^{(i)}_{\emptyset,\tilde{\bm{x}}}=2x_{2}-4.$
By combining the aforementioned basis functions, we obtain
$g^{(i)}_{{\bm{w}}_{4},\tilde{\bm{x}}}$
$\overset{\rm ii)}{=}$ $\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{{\bm{w}}_{4}})\!-\!\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{{\bm{w}}_{3}})\!-\!\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{{\bm{w}}_{2}})\!-\!\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{{\bm{w}}_{1}})$
$=$ $x_{1}x_{2}-2x_{1}-x_{2}+2.$
[…]
The set $\mathcal{G}_{i}$ can be viewed as the convex hull of the function space $\check{\mathcal{G}}_{i}$ spanned by a set of basis functions $\{g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}\}$. These basis functions are constructed based on the initial control law $\check{g}_{i}$, with respect to different subsets of the state variables $\bm{w}$} and anchor points $\tilde{\bm{x}}$. Each $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}$ captures the local variation characteristics of the control law in the corresponding subspace of the state, and they constitute a typical set of bases in the function space. By weighting and combining these basis functions, one can flexibly approximate or represent control strategies in the neighborhood of the initial control law.
@@
In this work, we assume that the $\mathcal{L}_{2}$-norm of the initial control law $\check{g}_{i}$ is bounded, i.e., $\|\check{g}_{i}\|_{\mathcal{L}_{2}}\leq M_{g}$. Consequently, the $\mathcal{L}_{2}$-norm of the basis functions $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}$ are also bounded, i.e., $\|g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}\|_{\mathcal{L}_{2}}\leq M_{g}$.
@@
For the considered system (1), there exists an optimal controller ${\bm{g}}^{*}=[g_{1}^{*},\ldots,g_{m}^{*}]^{\rm T},~g_{i}^{*}\in\mathcal{G}^{*}=\left\{g\mid\|g\|_{\mathcal{L}_{2}}\leq M_{g}\right\}$ that minimizes the cost function $J({\bm{g}})$.